### v2

# OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models

Abstract
Spatial reasoning is a key aspect of cognitive psychology and remains a bottleneck for current vision-language models (VLMs). While extensive research has aimed to evaluate or improve VLMs’ understanding of basic spatial relations, such as distinguishing left from right, near from far, and object counting, these tasks cover only the most elementary layer of spatial reasoning and are largely approaching saturation in the latest reasoning models. In this work, we introduce OmniSpatial, a comprehensive and challenging benchmark for spatial reasoning, grounded in cognitive psychology. OmniSpatial covers four major categories: dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking, with 50 fine-grained subcategories. Through careful manual annotation, we construct over 8.4K question-answer pairs. Extensive experiments show that both open- and closed-source VLMs exhibit significant limitations in comprehensive spatial reasoning. We also explore two strategies—PointGraph (explicit scene graph cues) and SpatialCoT (novel-view chain-of-thought)—to bolster spatial reasoning.
Overview of OmniSpatial Benchmark.
1 Introduction
Spatial reasoning plays a crucial role in bridging visual observation to robotic action (Huang et al., 2024b, Qi et al., 2024; 2025), autonomous driving, and AR/VR. For models to execute tasks effectively, they must understand spatial relationships to determine appropriate actions. To enhance the spatial understanding of Vision-Language Models, prior works (Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Ma et al., 2024, Song et al., 2024, Yuan et al., 2024, Yang et al., 2024, Qi et al., 2025) have integrated spatial information into datasets, enabling basic forms of spatial reasoning. Various benchmarks (Du et al., 2024, Szymanska et al., 2024, Shiri et al., 2024, Fu et al., 2024b, Song et al., 2024, Yang et al., 2024) have been introduced to systematically evaluate such capabilities, focusing on tasks like recognizing left and right, estimating depth, and constructing cognitive maps (Yang et al., 2024, Tolman, 1948, Momennejad et al., 2023). Additionally, spatial reasoning has been applied to manipulation tasks (Driess et al., 2023, Qi et al., 2025, Yuan et al., 2024), allowing systems to position objects according to specified spatial rules.
However, existing benchmarks still target basic spatial understanding, such as position relationship (left, right, front, back), proximity (near, far), and object counting. We use the latest reasoning models and agents to evaluate these benchmarks, such as o3 (OpenAI, 2025b) and Gemini-2.5-Pro (Reid et al., 2024). The results are shown in the lower left of OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models. These models have achieved over 90% accuracy on previous benchmarks such as SpatialBot-Bench (Cai et al., 2024) and EmbSpatial (Du et al., 2024), suggesting that these basic tasks are approaching saturation.
We believe complex spatial reasoning remains a significant challenge (Gardner, 2011, Baddeley, 1998, Previc, 1998, Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025). Human interaction with the physical world often involves interpreting ambiguous, dynamic, and context-dependent spatial relationships (Bar-Anan et al., 2006, Trope & Liberman, 2010, Ramalho et al., 2018). For example, in an emergency, knowing that an AED is “to the right of the door” is insufficient without understanding schematic diagrams, correlating maps with real-world environments, and planning an efficient route. Similarly, tasks like inserting a knife into a rack or flattening a box demand reasoning about object rotation, deformation, and spatial compatibility—far beyond static object placement.
From the perspective of cognitive psychology, complex spatial reasoning goes beyond simple relational judgments, encompassing dynamic world-knowledge reasoning, interactive spatial behavior with environments or agents, logical analysis of three-dimensional structures, and perspective-taking abilities. Motivated by these challenges, we introduce OmniSpatial, a comprehensive benchmark designed to capture the breadth and depth of spatial cognition. OmniSpatial systematically categorizes spatial reasoning into four core dimensions—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—thus providing a principled foundation for developing next-generation spatially- and physically-aware AI systems.
Table 1: Comparison with other spatial reasoning benchmarks. A comparison between OmniSpatial and other existing spatial reasoning benchmarks. OmniSpatial avoids template-based annotations, features highly diverse data, and includes a significantly larger number of tasks.
Dataset Embodied Task Categories Data Domain Data Annotation Data Scale Spatial QAs
EmbSpatial-Bench (Du et al., 2024) 6 Indoor (ScanNet, etc.) Template 2.2K 3.6K
Space3D-Bench (Szymanska et al., 2024) 6 Indoor (Replica) Manual 211 1K
Visual Spatial (Liu et al., 2023a) 7 MSCOCO Template 10K 10K
SpatialRGPT-Bench (Cheng et al., 2024) 12 Urban, Indoor, Sim Template 1.4K 1.4K
What’s up (Kamath et al., 2023) 6 Household, GQA, COCO Template 5K 5K
Spatial-MM (Shiri et al., 2024) 4 Internet Template 2.3K 2.3K
RoboSpatial (Song et al., 2024) 4 Indoor, tabletop Template 1M 3M
SpatialVLM (Chen et al., 2024) 2 WebLi Template 546 546
SpatialBot-Bench (Cai et al., 2024) 5 COCO,VG,RTX Manual 200 360
VSI-Bench (Yang et al., 2024) 8 Indoor Template 288 5K
OmniSpatial (Ours) 50 Internet Manual 6.5K 8.4K
The OmniSpatial benchmark includes images or video frames across diverse scenes, resolutions, lighting conditions, and weather patterns, collected from multiple countries across different continents. We evaluate state-of-the-art VLMs on our benchmark. Our findings indicate that, while current models perform well on conventional benchmarks, OmniSpatial presents a significantly greater challenge due to its comprehensive and complex task design.
Our key contributions are as follows:
  • We categorize visual-spatial reasoning into four key dimensions—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—broadening the scope of evaluation and guiding future research on spatial cognition in embodied & physics intelligence.
  • We develop the OmniSpatial dataset, which offers a diverse and challenging set of spatial tasks, serving as a comprehensive benchmark for assessing VLMs’ spatial reasoning capabilities.
  • We explore the enhancement of spatial reasoning in VLMs by incorporating auxiliary models in a chain-of-thought manner, demonstrating improved reasoning performance through this approach.
Figure 1: Benchmark Statistics of OmniSpatial: The distribution of tasks across 4 main categories.
2 Preliminaries: Visual–Spatial Reasoning
Spatial reasoning constitutes the cognitive bridge between visual perception and geometric understanding. We define visual–spatial reasoning as the capacity of an artificial system to infer, predict, and reason spatial properties of the world from visual observations. Formally, let an RGB observation stream be $\mathbf{I}_{1:T}$ and a task‑specific query be $q$. A model possesses visual–spatial reasoning that learn a mapping:
$f:(\mathbf{I}_{1:T},\,q)\;\longrightarrow\;a,$ (1)
where $a$ belongs to a well‑defined action or answer space whose correctness can be verified in the physical or simulated environment. This definition excludes non‑visual priors so that improvements can be attributed to visual reasoning itself, yet it remains compatible with multi‑modal extensions discussed in §1.
2.1 Taxonomy of Visual–Spatial Reasoning
Our taxonomy is motivated by two complementary perspectives. (i) Cognitive psychology foundations. Prior research on spatial cognition highlights partly independent faculties—such as visualization, mental rotation, perspective taking, and spatial updating—that can be systematically assessed in humans (Chabris et al., 2006, Meneghetti et al., 2022). These constructs provide a principled scaffold for analyzing how agents perceive, reason about, and act within space. (ii) Moving beyond basic spatial relations. Existing benchmarks are nearly saturated on simple tasks like left–right discrimination, front–back identification, and object counting (Chen et al., 2024, Cai et al., 2024, Du et al., 2024, Szymanska et al., 2024). Yet real-world embodied tasks demand richer reasoning (Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025, Stogiannidis et al., 2025, Lee et al., 2025) about scene dynamics, multi-step logic, physical interaction, and viewpoint transformation.
Guided by these considerations, we partition visual–spatial reasoning into four complementary dimensions: dynamic reasoning, complex spatial logic, spatial interaction, and perspective taking. Each dimension corresponds to a specific cognitive faculty and targets under-explored challenges in prior work, enabling us to probe a broader spectrum of spatial cognition while remaining grounded in psychological theory.
Dynamic Reasoning concerns inferring motion and temporal change from visual evidence. While our benchmark primarily uses static or sparsely sampled frames, such inference is crucial for adaptive decision-making in domains like robotics and navigation.
Complex Spatial Logic involves higher-order reasoning about relations, transformations, and geometric structures. It underpins problem-solving in design, engineering, and manipulation, where anticipating structural or relational changes is essential.
Spatial Interaction emphasizes reasoning guided by environmental constraints and task goals, covering skills such as path planning, obstacle avoidance, and context-aware action selection.
Perspective Taking captures the ability to adopt alternative viewpoints, supporting navigation, social cognition, and multi-agent coordination. It enables understanding relations from diverse perspectives and fosters flexible problem-solving.
2.2 Rationale for Classification
This taxonomy balances theoretical comprehensiveness with practical applicability. Dynamic reasoning highlights motion inference, complex logic captures abstract transformations, spatial interaction addresses real-time engagement, and perspective taking reflects cognitive flexibility. Together, these dimensions provide a framework for evaluating spatial reasoning in AI, robotics, and human cognition.
Figure 2: Tasks Demonstration of OmniSpatial. Several representative subtasks are selected for demonstration in each of the four task categories. Note: The questions above are slightly simplified for clarity and conciseness.
3 OmniSpatial: Comprehensive Spatial Reasoning Benchmark
3.1 Overview
We present OmniSpatial, a comprehensive benchmark designed to evaluate vision-language models on spatial reasoning. Rather than pursuing sheer data volume, OmniSpatial emphasizes diversity, structure, and rigor. It now consists of 8.4K carefully curated question–answer pairs, substantially larger than earlier prototypes, and covers a broad spectrum of scenarios that demand reasoning beyond pattern recognition.
The dataset integrates heterogeneous sources—web imagery, standardized cognitive tests, driving-exam questions, and prior dataset images such as MME (Liang et al., 2024) and HOI4D (Liu et al., 2022). This mixture enriches both realism and complexity: natural images capture everyday environments and architectures; psychology-inspired tasks introduce scientifically grounded challenges; driving exams provide safety-critical dynamic reasoning; and embodied datasets contribute varied resolutions, viewpoints, and human–object interactions.
Tasks are organized into 4 major categories and 50 fine-grained subtypes as shown in Figs.˜6 and 2, spanning from basic perspective-taking to dynamic motion prediction and spatial interaction in cluttered scenes. Each item is manually designed and reviewed through multi-round annotation, ensuring accuracy, consistency, and minimal ambiguity. This careful curation yields a benchmark that not only broadens the coverage of spatial reasoning but also establishes a reliable ground for evaluating and advancing future multimodal intelligence.
Figure 3: Data Construction of OmniSpatial.The pipeline collects images from multiple sources and ensures their quality and relevance through manual selection. Precise annotations are then applied, ensuring each question has a clear, unique answer while maintaining a natural, conversational expression. This process supports the effective training of VLMs in spatial reasoning tasks.
3.2 Benchmark Construction
3.2.1 Data Collection
As described in Section˜2, we define four spatial categories and corresponding task types. To build a diverse and information-rich dataset, we design targeted search strategies for each type, optimizing for relevance, diversity, and complexity.
Web Images. Web images form a major part of our data, especially for perspective-taking, dynamic reasoning, and spatial interaction. We use task-specific search terms (e.g., “indoor layout,” “furniture arrangement”) and append filters (“-ai,” “-generated”) to reduce synthetic content. Images are retrieved via Google’s Custom Search JSON API, Web RPA, and manual search, followed by strict filtering to remove irrelevant, low-resolution, or spatially trivial cases (e.g., isolated static objects). The resulting set balances realism and complexity, ensuring broad task coverage. All images are under MIT or CC-BY 4.0 license.
Exam-Based Test Questions. To capture abstract spatial logic, such as 3D transformations, rotations, and perspective shifts, we collect public spatial cognition tests through web scraping and manual curation. We categorize questions by focus and difficulty to maintain balance, removing redundant or knowledge-heavy items in favor of those targeting pure spatial reasoning. This refinement increases challenge diversity and improves benchmark quality.
Driving Test Questions. To evaluate reasoning in dynamic environments, we source tasks from three channels: (i) image-based multiple-choice questions from driving exam websites across at least three countries, (ii) online banks of standardized tasks like turning, lane changing, and parking, and (iii) interactive U.S. driving test videos, from which we extract frames, annotate bounding boxes, and design contextual queries (e.g., “Which bounding box indicates a potential traffic hazard?”). This combination yields realistic and challenging traffic scenarios, enhancing VLM adaptability to safety-critical reasoning tasks.
Existing Dataset Images  We integrate two key data sources: MME (Liang et al., 2024) and HOI4D (Liu et al., 2022). MME provides RGB-D data, allowing depth-based spatial inference. We leverage its depth information and manually propose physics-based questions such as “If a red car passes me in 5 seconds, what speed should I maintain?” This ensures realistic distance and motion-based reasoning. HOI4D contains extensive human-object interaction videos. We extract sequential frames to create motion prediction tasks, such as “Where will the hand holding the kettle move next?” By incorporating these datasets, we introduce real-world motion and interaction complexities, further strengthening VLMs’ dynamic reasoning capabilities. Furthermore, to extend our dataset to model training, we partition the dataset into a 1.5K test set and a 6.9K training set. The test set is entirely human-annotated and further integrates samples from several existing datasets. including SpatialViz  (Wang et al., 2025), PhysBench (Chow et al., 2025), ViewSpatial (Li et al., 2025), and DrivingVQA (Corbière et al., 2025), which substantially enhance the diversity of our dataset.
Figure 4: PointGraph: Enhance Spatial Reasoning through Additional Scene Graphs.
Figure 5: SpatialCoT: Enhancing Spatial Imagination through Novel View Synthesis.
3.2.2 Question-Answer Annotation
We design multiple-choice questions, including binary (true/false) and four-option formats, to enable standardized evaluation while minimizing annotation bias. To ensure naturalness, questions are phrased in conversational and context-rich styles (e.g., “If you are entering the classroom, on which side are the students?”) rather than rigid templates (e.g., “Is [A] on the right of [B]?”). This encourages models to rely on contextual and relational reasoning rather than memorized patterns.
To guarantee clarity and answer uniqueness, we perform multiple rounds of validation and resolve ambiguous references through cross-checking among annotators. Six trained annotators—initially the authors—were involved in labeling. Each annotator underwent task-specific training, and annotations were cross-validated to reduce subjectivity. The inter-annotator agreement reached Krippendorff’s $\alpha=0.84$, indicating a high level of consistency.
3.3 Improving Visual Spatial Reasoning Abilities
3.3.1 PointGraph: Explicit Modeling of Object Relationships
To strengthen models’ understanding of spatial relations, we introduce PointGraph, which constructs structured representations of object relationships within an image. Concretely, we employ open-vocabulary grounding models such as Florence-2 to localize multiple objects and extract their centers and bounding boxes. These detections are then assembled into a JSON-style scene graph encoding object identities and relative positions. By concatenating this structured spatial description with the original query, VLMs are provided with explicit geometric cues that facilitate more accurate reasoning about distances, directions, and configurations (see Fig.˜5).
3.3.2 SpatialCoT: Stimulating Spatial Imagination via Novel Views
Human spatial reasoning often relies on mental imagery—the ability to imagine how a scene looks from different viewpoints. Inspired by this, we design SpatialCoT, which augments visual inputs with 3D novel views to enrich spatial imagination (Shi et al., 2023, Qi et al., 2023b, Liu et al., 2024d). Specifically, we adopt InstantMesh (Xu et al., 2024a) to synthesize six additional perspectives for each input image and compose them into a multi-view collage. This collage is then fed, along with the question, into the VLM as part of a chain-of-thought prompting pipeline. The additional perspectives provide strong geometric priors, helping models disambiguate occlusions, perspective-taking, and other view-dependent reasoning tasks.
Table 2: Evaluation on OmniSpatial-test. All models were tested 5 times and averaged to reduce randomness. Dark green indicates the best result and light green indicates the second-best result within the group. Gemini-2.5-Pro (Reid et al., 2024), InternVL3-78B (Zhu et al., 2025), and SoFar (Qi et al., 2025) achieve the best performance within their respective groups.
Method
Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
Manipulate
Motion
Analysis
Traffic
Analysis
Locate
Geospatial
Strategy
Pattern
Recognition
Geometric
Reasoning
Ego
Centric
Allo
Centric
Hypothetical
Blind Evaluation
Random Choice 24.98 - 24.86 26.30 25.88 23.43 27.27 21.44 24.77 22.55 24.84 25.78
GPT-3.5-turbo (Roumeliotis & Tselikas, 2023) 30.67 - 38.38 29.19 38.35 28.76 36.91 0.82 24.00 42.16 33.67 35.90
GPT-4-turbo (OpenAI, 2023) 34.06 - 42.97 37.40 41.18 28.95 40.00 22.27 26.32 31.37 33.99 35.42
Proprietary Models
GPT-4o-mini-2024-07-18 (Hurst et al., 2024) 42.64 8 55.95 50.29 54.59 43.43 44.91 22.47 29.42 61.57 36.76 34.22
GPT-4o-2024-11-20 (Hurst et al., 2024) 47.81 5 65.54 57.23 56.47 52.38 54.09 26.29 25.48 75.98 39.49 39.76
GPT-4.1-nano-2025-04-14 (OpenAI, 2025a) 42.62 9 50.90 53.85 54.90 40.95 42.42 24.40 30.11 53.59 37.23 33.73
GPT-4.1-mini-2025-04-14 (OpenAI, 2025a) 48.87 4 64.32 56.53 59.06 60.19 56.36 29.28 30.19 72.55 39.57 39.28
GPT-4.1-2025-04-14 (OpenAI, 2025a) 51.78 2 66.22 64.74 60.00 65.33 60.18 31.75 30.06 70.98 40.64 39.04
Claude-3-5-sonnet-20241022 (Anthropic, 2024) 46.86 6 54.05 54.57 58.12 68.38 53.09 26.60 31.74 70.00 34.79 39.52
Claude-3-7-sonnet-20250219 (Anthropic, 2024) 47.53 5 57.57 55.95 56.71 63.81 59.09 29.48 28.39 72.16 36.06 36.63
Gemini-2.0-flash-lite-02-05 (Anil et al., 2023) 44.03 8 59.19 46.71 60.24 49.52 53.27 21.65 31.23 66.47 36.81 38.80
Gemini-2.0-flash-exp (Anil et al., 2023) 48.40 2 61.89 56.01 51.76 63.43 59.09 20.82 33.81 72.75 39.20 39.28
Gemini-2.5-flash-preview-05-20 (Anil et al., 2023) 52.12 1 67.57 62.72 68.24 73.33 60.91 38.14 34.19 75.49 35.90 33.73
Reasoning Models
o1-2024-12-17 (Jaech et al., 2024) 50.36 6 71.62 60.98 57.65 63.81 60.00 39.18 27.10 71.57 38.03 36.14
o4-mini-2025-04-16 (OpenAI, 2025b) 52.77 3 72.97 59.83 60.00 73.33 61.82 34.02 36.77 73.53 40.69 40.96
o3-2025-04-16 (OpenAI, 2025b) 56.33 1 71.89 66.18 61.18 68.57 65.45 40.21 29.68 77.06 48.40 48.19
Claude-3-7-sonnet-20250219-thinking (Anthropic, 2024) 48.62 7 57.21 59.73 53.73 67.94 57.27 30.24 28.17 68.63 37.94 36.95
Gemini-2.5-flash-05-20-thinking (Anil et al., 2023) 53.16 3 70.27 64.74 61.18 72.38 58.18 35.05 36.13 74.12 40.96 32.53
Gemini-2.5-pro-preview-05-06 (Anil et al., 2023) 55.19 2 67.57 71.39 62.35 75.24 64.55 43.30 34.84 74.51 38.03 37.35
Open-source Models
LLavA-1.5-vicuna-7B (Liu et al., 2024c) 34.97 15 54.46 31.23 35.29 36.19 33.94 29.01 24.18 55.60 34.66 36.14
LLaVA-onevision-qwen2-7B (Li et al., 2024a) 35.68 14 43.24 38.15 32.94 29.52 41.82 28.87 22.58 47.06 36.17 37.35
LLaVA-onevision-qwen2-72B (Li et al., 2024a) 45.66 6 62.16 50.29 54.12 60.95 56.36 22.68 25.81 76.47 37.23 33.73
Gemma-3-4B (Kamath et al., 2025) 39.79 11 41.89 49.71 56.47 27.62 36.36 23.71 24.52 59.80 36.17 38.55
Gemma-3-12B (Kamath et al., 2025) 43.71 8 54.05 54.91 54.12 47.62 45.45 16.49 30.32 63.73 36.70 33.73
Gemma-3-27B (Kamath et al., 2025) 44.75 7 56.76 55.78 57.65 50.48 52.73 27.84 29.03 64.71 33.51 32.53
InternVL3-2B (Zhu et al., 2025) 37.98 13 50.00 40.58 43.29 40.00 40.55 21.86 28.52 55.49 35.11 33.01
InternVL3-8B (Zhu et al., 2025) 41.60 9 52.43 40.87 48.94 51.05 44.77 24.95 28.63 64.20 38.62 40.96
InternVL3-14B (Zhu et al., 2025) 45.94 5 54.32 60.17 50.35 51.81 51.45 28.04 28.26 68.04 35.37 34.46
InternVL3-38B (Zhu et al., 2025) 48.48 2 63.42 63.58 54.59 58.29 50.55 29.90 28.52 72.16 36.76 33.49
InternVL3-78B (Zhu et al., 2025) 49.33 1 63.78 63.12 56.24 59.24 51.45 27.63 30.19 74.51 38.46 35.90
Qwen-VL2.5-3B (Wang et al., 2024b) 40.30 10 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
Qwen-VL2.5-7B (Wang et al., 2024b) 39.18 12 58.38 35.09 50.12 45.33 44.00 31.13 29.42 64.51 33.19 37.35
Qwen-VL2.5-32B (Wang et al., 2024b) 47.36 4 63.06 55.09 51.76 66.29 56.91 26.39 27.48 68.04 37.50 40.24
Qwen-VL2.5-72B (Wang et al., 2024b) 47.85 3 58.38 60.12 50.12 59.81 53.64 26.19 33.03 71.37 36.81 36.39
Specialized Spatial Reasoning Models
SpaceMantis-13B (Chen et al., 2024) 36.36 6 47.03 36.59 40.94 34.86 33.09 22.27 24.39 49.22 38.25 39.28
SpaceQwen2.5-VL-3B (Chen et al., 2024) 40.25 3 58.11 39.88 41.18 40.95 40.91 29.90 25.81 63.73 38.83 39.76
SpaceThinker-Qwen2.5VL-3B (Chen et al., 2024) 40.42 2 47.84 53.06 43.29 35.43 38.73 24.33 28.00 58.04 35.11 31.08
SpatialBot-3B (Cai et al., 2024) 35.68 6 43.24 38.15 32.94 29.52 41.82 28.87 22.58 47.06 36.17 37.35
RoboPoint-vicuna-v1.5-7B-lora (Yuan et al., 2024) 35.85 6 57.03 28.61 34.82 37.33 40.55 29.90 22.71 50.20 38.72 40.96
RoboPoint-vicuna-v1.5-13B (Yuan et al., 2024) 34.60 5 55.68 28.15 42.82 32.19 32.55 24.12 27.74 49.02 37.66 33.49
SoFar-Qwen2.5VL-3B (Qi et al., 2025) 45.14 1 56.49 51.16 54.12 53.14 52.73 31.75 22.88 71.60 36.56 41.69
Human Evaluation
Human 92.63 - 94.62 96.07 91.38 95.11 92.15 89.02 85.90 98.53 94.30 90.26
4 Experiments
4.1 Evaluation Setup
To systematically assess spatial reasoning, we evaluate models on the OmniSpatial benchmark under a unified protocol. Our evaluation covers both proprietary and open-source systems, spanning general-purpose VLMs, reasoning-oriented LLMs, and spatially specialized models, thereby ensuring broad coverage and fair comparison. We consider four groups of state-of-the-art models:
  • Proprietary Models. These include the GPT-4o and GPT-4.1 families (Hurst et al., 2024), the Claude series (Anthropic, 2024), and Gemini series (Anil et al., 2023). All are accessed through APIs under zero-shot settings with standardized system prompts (see Appendix˜D).
  • Reasoning Models. We categorize models that explicitly employ long chain-of-thought reasoning, often enhanced through reinforcement learning, such as o1, o3, o4-mini (Jaech et al., 2024, OpenAI, 2025b), and Gemini-2.5 (Anil et al., 2023, Reid et al., 2024). Because their outputs are less amenable to strict parsing, we additionally rely on an automatic judge to compare answers against ground truth, following the LLM-as-a-Judge paradigm (Zheng et al., 2023).
  • Open-Source Models. This set includes Qwen-VL (Bai et al., 2023), InternVL (Zhu et al., 2025), Gemma (Kamath et al., 2025), and LLaVA-OneVision (Liu et al., 2023b). These are locally deployed with standardized prompts to ensure reproducibility.
  • Specialized Spatial Models. We further benchmark models designed specifically for spatial reasoning, including SpatialVLM (Chen et al., 2024), RoboPoint (Yuan et al., 2024), SpatialBot (Cai et al., 2024), and SoFar (Qi et al., 2025). These systems incorporate explicit spatial signals such as metric 3D information, point affordances, or semantic orientation to improve reasoning.
4.2 Evaluation Metrics
We measure accuracy on multiple-choice questions. For standard proprietary and open-source models, we test four output protocols: direct answer, regular-expression parsing, JSON parsing, and LLM-as-a-Judge (Zheng et al., 2023). For reasoning-oriented models with unstructured CoT outputs, correctness is assessed by GPT-4.1-mini against ground truth. Ablations of these evaluation strategies are provided in Section˜C.1.
Table 3: Camparison of textual Chain-of-Thought and PointGraph on OmniSpatial-test.
Method
Avg. Improve Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
Manipulate
Motion
Analysis
Traffic
Analysis
Locate
Geospatial
Strategy
Pattern
Recognition
Geometric
Reasoning
Ego
Centric
Allo
Centric
Hypothetical
GPT-4.1-mini - - - - - - - - - - - -
(w/o CoT) 48.86 - 64.05 58.55 57.65 59.43 56.91 28.87 34.06 68.82 37.18 41.20
(w/ Zero-shot CoT) 49.81 +0.95 62.97 58.96 59.06 62.48 58.55 27.63 32.52 69.41 40.11 40.96
(w/ Manual CoT) 49.76 +0.90 65.68 58.90 58.59 64.38 56.91 28.45 32.13 69.61 39.31 41.20
(w/ PointGraph) 50.49 +1.63 67.57 62.14 57.65 64.76 58.18 28.87 30.32 70.59 38.83 42.17
Gemini-2.5-Flash - - - - - - - - - - - -
(w/o CoT) 51.47 - 66.22 65.90 63.53 71.43 66.36 32.99 34.84 70.59 31.91 38.55
(w/ Zero-shot CoT) 51.53 +0.06 63.51 61.27 58.82 67.62 65.45 42.27 34.84 79.41 35.90 32.53
(w/ Manual CoT) 52.12 +0.65 67.57 62.72 68.24 73.33 60.91 38.14 34.19 75.49 35.90 33.73
(w/ PointGraph) 53.23 +1.76 62.16 69.94 64.71 67.62 59.09 29.90 38.06 74.51 37.77 37.35
Qwen-VL2.5-3B - - - - - - - - - - - -
(w/o CoT) 41.45 - 58.65 43.06 39.53 50.67 48.73 32.78 22.58 61.96 37.66 37.35
(w/ Zero-shot CoT) 40.64 -0.81 59.73 43.87 46.12 48.38 43.27 25.36 22.84 59.61 36.54 37.59
(w/ Manual CoT) 40.07 -1.38 55.68 46.65 47.29 40.57 46.00 28.04 24.39 60.39 33.35 31.57
(w/ PointGraph) 44.36 +2.91 55.68 55.20 48.94 52.19 52.36 29.90 25.55 66.08 35.11 31.08
4.3 Main Results
Overall Model Performance  As illustrated in Table˜2, we observe the following findings: (i) Proprietary reasoning models, such as ChatGPT o3 (OpenAI, 2025b) and Gemini-2.5-pro (Anil et al., 2023), achieve the highest performance, surpassing a  56% overall success rate; however, there remains a significant gap compared to human-level understanding, and they require a lot of inference time and tokens. (ii) Open-source models also demonstrate competitive results, with large-scale models like InternVL3-78B (Zhu et al., 2025) and Qwen-VL2.5-72B (Wang et al., 2024b) achieving comparable performance to GPT-4.1-mini and Gemini-2.0-flash-exp. (iii) Specialized Spatial Reasoning Models, due to limitations in dataset coverage and model capacity, struggle to achieve substantial improvements on comprehensive benchmarks.
Category-wise Analysis  We observe notable performance differences across spatial reasoning categories: (i) Leveraging their extensive world knowledge and local understanding capabilities, proprietary models have demonstrated strong performance in Dynamic Reasoning and Spatial Interaction, indicating that reasoning models possess high proficiency in temporal understanding, spatial relationship analysis, and map-based comprehension. (ii) For Pattern Geometric Reasoning, which involves spatial imagination in planar geometry, even reasoning models designed for extended thinking can only achieve an accuracy of around 30% to 40%, slightly surpassing the random baseline. (iii) Current models exhibit limited perspective-taking abilities, predominantly analyzing scenarios from an ego-centric viewpoint while struggling to imagine perspectives from others’ viewpoints.
Table 4: Performance of Spatial CoT on OmniSpatial Perspective-Taking track.
Method Avg. Improve
Ego
Centric
Allo
Centric
Hypothetical
GPT-4.1-mini
(w/ Zero‑shot CoT) 45.56 - 69.41 40.11 40.96
(w/ Spatial CoT) 47.58 +2.02 69.43 42.37 44.34
Qwen‑VL2.5‑3B
(w/ Zero‑shot CoT) 40.89 - 59.61 36.54 37.59
(w/ Spatial CoT) 42.90 +2.01 60.80 39.25 37.44
Impact of PointGraph & Spatial CoT  To test whether structured segmentation improves performance, we apply PointGraph as a preprocessing step for GPT-4.1, Gemini-2.5-flash and Qwen-VL2.5-7B. Results in Table˜3 show a clear accuracy boost, particularly in the Dynamic Reasoning and Perspective-Taking Track, validating the benefits of integrating structured object representation, while traditional textual CoT difficult to bring about significant improvement. Fig.˜5 further demonstrates the effectiveness of our proposed Spatial CoT on the OmniSpatial Perspective-Taking track. Through novel view synthesis facilitated by InstantMesh, both GPT-4.1 and Qwen-VL2.5-7B exhibit significant performance improvements, validating the effectiveness of explicit spatial imagination.
4.4 Training Exploration
Supervised Training on OmniSpatial-train.  We further study whether OmniSpatial-train can effectively teach spatial skills instead of merely fitting templates. Starting from a strong open-source baseline, supervised fine-tuning on 6.9K samples yields a substantial +7.82 point average gain over zero-shot, with consistent improvements across dynamic, interaction, and perspective-taking oriented tracks. In contrast, training on a much larger 200K template-style corpus that follows the construction process of VSI-Bench (Yang et al., 2024) brings only a marginal +1.29 average gain, underscoring the value of diverse, manually curated spatial tasks over synthetic templates.
Table 5: Training Exploration on OmniSpatial-test.
Method Avg. Improve
Manipulate
Motion
Analysis
Traffic
Analysis
Locate
Geospatial
Strategy
Pattern
Recognition
Geometric
Reasoning
Ego
Centric
Allo
Centric
Hypothetical
Qwen-VL2.5-3B - - - - - - - - - - - -
Zero-shot 40.30 - 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
+ OmniSpatial-train (6.9K) 48.12 +7.82 59.19 51.16 50.12 51.81 45.45 34.02 31.74 68.63 36.76 35.90
+ Template corpus (200K) 41.59 +1.29 50.90 46.71 40.94 45.33 40.55 28.87 24.39 72.55 35.11 33.01
Generalization to VSI-Bench.  To examine cross-benchmark generalization, we adopt the SpaceR-7B pipeline and compare supervised training with and without OmniSpatial. Adding OmniSpatial improves the overall score on VSI-Bench from 41.68 to 43.68. Notably, it boosts categories requiring ordering, counting, and metric/room-size reasoning (e.g., appearance_order, obj_counting, obj/room_size). These results indicate that OmniSpatial provides complementary supervision that transfers to external spatial tasks rather than overfitting to in-benchmark patterns.
Table 6: Generalization on VSI-Bench. Training with OmniSpatial yields consistent gains.
Method overall
Appearance
Order
Obj Abs
Distance
Obj
Counting
Obj Rel
Distance
Obj
Size
Room
Size
Route
Planning
Obj Rel
Direction
Qwen-VL2.5-3B - - - - - - - - -
Zero-shot 34.06 34.63 11.39 40.18 36.20 46.11 38.58 31.96 36.92
+ SpaceR-7B 41.68 46.60 19.63 55.27 34.23 58.32 35.17 34.54 42.84
+ OmniSpatial 43.68 58.25 15.13 57.36 34.37 60.99 41.11 34.02 44.16
5 Related Works
5.1 Benchmarking Spatial Reasoning
Various studies have introduced innovative benchmarking methodologies (Szymanska et al., 2024, Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Song et al., 2024, Qi et al., 2025, Ray et al., 2024) to advance spatial reasoning evaluation. Spatial VQA (Du et al., 2024) was among the first to incorporate spatial information into vision-language models, enabling fundamental spatial relationship reasoning. SpatialBot (Cai et al., 2024) categorized spatial reasoning into various hierarchical levels, extending its applicability to robotic manipulation tasks. RoboSpatial (Song et al., 2024) proposes a large-scale template-based spatial relationship dataset, focusing on positional relationships from different perspectives. VSI-Bench (Yang et al., 2024) combined video data (Chandrasegaran et al., 2024, Liu et al., 2023c, Fu et al., 2024a, Li et al., 2024b, Fang et al., 2024b) with cognitive maps (Momennejad et al., 2023, Apostolopoulos & Groumpos, 2023) to simulate human-like spatial cognition and optimize reasoning in dynamic environments. Recently, SoFar (Qi et al., 2025) introduced the 6-DoF SpatialBench to evaluate the understanding of orientation.
While these benchmarks have made significant contributions, a unified framework encompassing a wide range of complex spatial reasoning tasks remains lacking. Inspired by prior spatial reasoning research (Xu et al., 2024b, Wang et al., 2024a, Lin et al., 2014, Nwankwo et al., 2024), we identified several limitations in existing benchmarks, such as reliance on generated images (Szymanska et al., 2024, Fu et al., 2024b, Kamath et al., 2023, Liu et al., 2023a, Rajabi & Kosecka, 2023, Shiri et al., 2024), LLM-generated templates (Linghu et al., 2024, Du et al., 2024, Rädsch et al., 2025), and domain-specific focus (Xie et al., 2025, Chow et al., 2025, Danish et al., 2024). These issues hinder their comprehensiveness and real-world applicability. To address these gaps, our study proposes a comprehensive and integrative spatial reasoning benchmark.
6 Conclusion
We introduce OmniSpatial, a benchmark for comprehensive visual–spatial reasoning. OmniSpatial distills spatial cognition into four primary categories—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—spanning 50 fine-grained subtasks and 8.4 K manually-curated question–answer pairs. Extensive experiments show that state-of-the-art proprietary and open-source VLMs peak at 57% accuracy—over 30 points below human performance—struggling especially with geometric reasoning and non-egocentric perspective taking. To bridge these gaps, we introduce PointGraph for structured scene-graph reasoning and SpatialCoT for viewpoint-aware chain-of-thought, both yielding consistent gains and underscoring the value of structured and multi-view reasoning.
Appendix A Detailed Task Design
OmniSpatial aims to comprehensively evaluate the spatial reasoning capabilities of Vision-Language Models, covering four major categories: Dynamic Reasoning, Complex Logic, Spatial Interaction, and Perspective Taking. Each category not only focuses on different types of spatial reasoning tasks but also includes challenges based on real-world application scenarios. This approach helps researchers better understand and enhance models’ multi-domain spatial reasoning abilities. The following presents the underlying considerations and practical value behind each task design.
A.1 Dynamic Reasoning
The Dynamic Reasoning category focuses on the model’s understanding of object movement and its changes, assessing the ability to make accurate judgments in uncertain or rapidly changing environments. Spatial dynamics are critical not only in robot control but also have broad applications in fields like autonomous driving and intelligent surveillance.
A.1.1 Manipulation
Operational Position Selection  This task evaluates how models determine the optimal interaction point with objects in complex environments. Selecting the best grasping point can prevent tilting or damage to objects and improve the efficiency and precision of operations. This task is crucial in robotic grasping, especially when environmental conditions are unstable.
Movement Direction Determination  This task assesses the model’s ability to predict the movement direction of an object or itself, providing decision support for automated systems and helping to optimize robot motion strategies.
Intent Recognition  Intent recognition involves inferring the purpose or goal behind a movement. This task is particularly important for contextual analysis, such as determining whether a person is reaching for a door handle to open or close a door. This capability enhances the reasoning of human-robot interactions and optimizes the interaction experience of intelligent assistants and robots.
A.1.2 Motion Analysis
Uniform Motion  The model’s ability to reason about uniform motion reflects its fundamental understanding of time and spatial relationships, such as estimating the speed or time required for a target to move. This is applicable in areas like object tracking and path prediction, such as estimating vehicle travel time or train arrival time.
Variable Motion  Variable motion analysis focuses on understanding acceleration and deceleration processes. By predicting the position changes of an object during variable motion, the model can better simulate dynamic phenomena in the physical world, such as calculating braking distance for vehicles. This is critical in autonomous driving and robot control.
Spatial Compatibility  Tasks that assess whether an object fits within a specific space directly relate to the precision of robots and automated devices in real-world operations. For instance, determining whether luggage can fit into an overhead compartment is applicable in logistics and automated warehouses, helping systems make effective object adaptation decisions.
A.2 Complex Logic
The Complex Logic category focuses on higher-order spatial reasoning, including tasks such as geometric transformations and pattern recognition, challenging the model’s ability to abstractly understand and reason in multi-dimensional spatial environments.
A.2.1 Pattern Recognition
Style  Style recognition tasks assess the model’s ability to infer visual rules in structured patterns. Typical operations include completing missing parts, combining or subtracting shapes, comparing similarities and differences, and performing visual logic like black-white inversion. This skill is crucial for tasks like intelligence tests and diagrammatic reasoning.
Quantity  Quantity-based tasks evaluate the model’s ability to perform visual numerosity reasoning, focusing on implicit quantitative patterns such as the number of points, lines, regions, elements, or strokes. These tasks require abstract counting under diverse spatial configurations, often without explicit numerical labels or symbols.
Attributes  Attribute-based tasks evaluate the model’s ability to reason about non-numeric visual properties such as symmetry (axial or radial), curvature, and openness. These tasks require the recognition of structural features that do not rely on quantity but rather on geometric or perceptual traits.
Location  Location tasks evaluate the model’s ability to reason about spatial changes such as translation, rotation, and reflection. The focus is on how objects shift in space while preserving their structure. This skill is essential for grid-based reasoning, motion prediction, and geometric pattern understanding.
A.2.2 Geometric Reasoning
Polyhedron Unfolding  This task examines whether the model can infer the 2D net of a 3D object, reflecting its ability to mentally construct spatial layouts. It is particularly relevant to applications such as packaging design, industrial manufacturing, and aerospace engineering.
Sections and Projections  By evaluating how models interpret 2D cross-sections or projections from different viewpoints, this task challenges their capacity to connect visual appearance with internal structure—key to fields like medical imaging, architecture, and mechanical design.
Mental Rotation  Mental rotation requires the model to simulate the rotation of objects in mind and track changes across views. It underpins spatial reasoning in tasks ranging from CAD modeling to object manipulation in virtual and augmented reality.
Assembly  This task involves reasoning about how separate parts fit together into a coherent whole, testing the model’s understanding of geometric constraints. It has broad relevance in robotics, structural analysis, and physical assembly planning.
Analytical Geometry  These tasks are inspired by classic geometry problems, requiring models to reason about spatial relations using angles, distances, and symmetry. They bridge mathematical logic with visual structure, supporting applications in structured reasoning and spatial abstraction.
A.3 Spatial Interaction
The Spatial Interaction category evaluates a model’s ability to reason about interactions with objects and environments. It includes tasks such as Traffic Analysis, Localization, and Geospatial Strategy, reflecting the model’s understanding and application of spatial knowledge in real-world scenarios.
A.3.1 Traffic Analysis
Anomaly Detection  This task focuses on identifying potential dangers or traffic violations in complex scenes, such as unsafe following distances or unusual vehicle behavior. It plays a key role in ensuring safety in autonomous driving systems.
Sign Recognition  This task evaluates the model’s ability to detect and interpret traffic signs, including speed limits, no-entry zones, and yield signs. Accurate recognition is critical for safe and rule-compliant decision-making.
Action Recognition  This task involves identifying or predicting the actions of traffic participants, such as driver gestures, police signals, or pedestrian intentions. It is important for understanding dynamic human behavior in traffic environments.
Risk Detection  This task aims to detect immediate hazards in the environment, such as an opening car door or a pedestrian crossing the road. Timely detection supports effective avoidance and control strategies.
Behavior Guidance  This task provides context-aware behavioral suggestions, such as advising to turn off high beams or reminding that parking is prohibited. It enhances overall driving safety and compliance.
Contextual Analysis  This task assesses the model’s ability to interpret spatial relations and behaviors based on environmental cues. For example, estimating wind conditions to anticipate overtaking behavior or understanding road status to infer potential risks.
A.3.2 Localization
UI Interaction  This task requires the model to determine which icon should be selected within a user interface based on contextual understanding, and to accurately localize its position. It reflects the model’s ability to integrate semantic interpretation with spatial reasoning, supporting applications in intelligent assistants and automated interface control.
Object Detection  This task involves identifying specific target objects within an image. It is often paired with spatial localization to jointly assess what the object is and where it is located.
Spatial Localization  This task focuses on determining the precise position of objects within a scene. It is commonly evaluated alongside object detection to answer questions like “What is at this location?” or “Where is this object?”
Pose Estimation  This task estimates the orientation and spatial configuration of objects, such as detecting whether a cup is upright or tipped over. It is frequently integrated with spatial localization to enable more nuanced scene understanding.
A.3.3 Geospatial Strategy
Spatial Recognition  Assesses the model’s capacity to identify spatial structures such as rooms, corridors, or zones within a scene. This is essential for semantic navigation and indoor mapping.
Location Recognition  This task evaluates the model’s ability to identify specific locations on a map or scene, such as recognizing the position marked as “You are here” or locating a designated landmark. It reflects the model’s capacity to associate spatial markers with real-world positions.
Region Recognition  Focuses on distinguishing and classifying regions in a broader spatial context, such as residential vs. industrial zones on a map.
Route Interpretation  Tests the model’s ability to follow or explain a route depicted in a map or scene. It requires understanding directional arrows, route labels, and spatial transitions.
Route Design  Involves selecting or generating an optimal path to reach a given goal, considering spatial constraints and possible alternatives.
Route Selection  Compares multiple candidate routes and chooses the most suitable one based on efficiency, safety, or contextual requirements.
Navigation  Evaluates the model’s ability to understand smartphone or in-vehicle navigation interfaces, including interpreting turn-by-turn directions, identifying route segments, and understanding map overlays. This is crucial for building intelligent voice assistants and real-time guidance systems.
Map/Scene Conversion  Tests the ability to mentally convert between map views and real-world scenes, which is critical in correlating schematic representations with physical surroundings.
Legend Recognition  It requires identifying and interpreting map symbols (e.g., stairs, elevators, emergency exits) and a foundational skill in navigation and spatial reasoning.
Terrain Identification  Focuses on distinguishing types of terrain (e.g., flat, uphill, water-crossing), which is essential for planning safe and feasible paths in outdoor navigation or robotics.
A.4 Perspective Taking
This category evaluates the model’s ability to understand spatial relationships from different viewpoints. Since changes in perspective directly affect what is observed, the ability to reason across varying angles is essential for robotic perception and interaction.
A.4.1 Egocentric
Count  Counting the number of visible objects from the current perspective is crucial for dynamic interaction. For example, in robotic grasping tasks, knowing how many targets are visible helps determine the appropriate operation strategy.
Size  Judging the size of an object from the observer’s viewpoint aids robots or virtual systems in depth perception, helping assess whether an object can be grasped or properly placed.
Direction  This refers to the direction directly seen by the observer. It is especially important in autonomous driving scenarios, where understanding object movement helps predict traffic conditions and enables timely responses.
Order  Analyzing the arrangement of multiple objects in an image is essential for robotic operations, helping prioritize which objects to interact with first.
Distance  The distance between objects as perceived from the observer’s viewpoint is a key capability in navigation systems, enabling path planning and obstacle avoidance.
A.4.2 Allocentric
Count  Understanding how the perceived quantity of objects changes under different viewpoints is key. Due to variations in position and orientation, occlusion often occurs—for example, a driver may have blind spots, while a road surveillance camera can see objects hidden in those areas. This task evaluates the model’s ability to judge the difference in object counts from various observation points.
Size  This task involves evaluating an object’s size from a specified observer’s viewpoint. Due to the general principle that closer objects appear larger and distant ones appear smaller, objects of the same size may look different to different observers. The model is expected to either infer the true size based on reference objects or estimate how perceived size changes with viewing position.
Direction  This task emphasizes judging directions from abstract positions, such as another agent’s viewpoint or a map. The answers often differ from what is directly observed, requiring one to adopt the target’s perspective—engaging in perspective-taking. It is crucial not only for large-scale path planning but also for understanding an object’s intrinsic orientation.
Order  This task requires observing the arrangement of objects from a specified viewpoint—for example, the seating order of students in the front row as seen by a teacher on the podium, which is exactly reversed from what a camera at the back of the classroom would capture. Only by understanding “what the target sees” can one make accurate predictions or judgments about the scene.
Distance  Differences in the observer’s position and orientation lead to variations in perceived size and distance. For example, in the same scene, a photographer taking pictures of the same object from different camera angles will capture different impressions of its size and distance. Changing the camera position essentially means changing the perspective. This ability is vital for coordination and environmental assessment in human-robot interaction or multi-robot collaborative tasks.
A.4.3 Hypothetical Perspective Taking
This category focuses on imagining the scene from a specified but non-existent viewpoint, requiring the model to mentally adopt a fictional position—an advanced form of perspective-taking in spatial reasoning.
Count  Predicting how many target objects would be visible from a hypothetical viewpoint—for example, a person standing at the opposite corner of the street may see a different distribution of objects. The model must reason about occlusion, orientation, and visibility from the imagined perspective.
Size  Inferring the apparent size of objects from a hypothetical location and direction. For instance, the same object may appear larger when viewed up close or smaller when viewed from above. The model needs to simulate how visual scale changes with altered viewpoints.
Direction  Reasoning about how an object’s direction appears from a location where no observer is present. For example, a pedestrian walking toward a doorway would appear “head-on” from the entrance, but present a different direction from a side view.
Order  Simulating the arrangement of objects from another location helps assess how spatial sequences change across viewpoints. For example, the seating order seen from the podium may be the reverse of what’s seen from the back of the room.
Distance  Estimating relative distances between objects from a hypothetical position requires mentally adopting a new viewpoint. This supports effective planning and coordination in tasks such as multi-robot navigation or collaborative manipulation.
Appendix B Additional Related Works
B.1 Spatial Vision-Language Models
Spatial vision-language models integrate computer vision (He et al., 2022, Oquab et al., 2024, Radford et al., 2021) and natural language processing (Bai et al., 2023, Wang et al., 2024b, Brown et al., 2020, Anil et al., 2023, Touvron et al., 2023a; b, Zhang et al., 2024b) to enhance machine understanding of spatial relationships. Recent research (Qi et al., 2024, Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Song et al., 2024, Yuan et al., 2024, Qi et al., 2025) has increasingly focused on extending vision-language models to support spatial reasoning in dynamic and 3D environments. One of the pioneering works in this domain, SpatialVLM (Chen et al., 2024), has significantly advanced spatial reasoning by constructing an RGB-D based visual question-answering dataset. These models (Hurst et al., 2024; 2024, Reid et al., 2024, Wang et al., 2024b) effectively process multimodal data containing spatial information, serving as a bridge between visual perception and linguistic reasoning.
Further advancements (Liu et al., 2024b; a, Ramakrishnan et al., 2024, Tang et al., 2024, Yang et al., 2023, Rozanova et al., 2021, Wu et al., 2024, Stogiannidis et al., 2025) include SpatialRGPT (Cheng et al., 2024), which extends RGB-D spatial understanding by incorporating 3D scene graphs and spatial data to improve inference capabilities. Similarly, SpatialBot (Cai et al., 2024) explores hierarchical deep reasoning mechanisms to handle depth and spatial structures in complex environments. Recently, SoFar (Qi et al., 2025) proposed semantic orientation and trained an open-world orientation model to enhance the orientation understanding of VLMs, significantly improving spatial understanding and robotic operation capabilities (Cho et al., 2024, Driess et al., 2023, Mu et al., 2023, Zitkovich et al., 2023, Wake et al., 2024, Huang et al., 2024b, Nasiriany et al., 2024, Duan et al., 2024, Huang et al., 2024a, Qi et al., 2025, He et al., 2025a, Zhang et al., 2025b). However, existing research remains limited in addressing the full complexity and comprehensiveness of spatial reasoning tasks. To bridge this gap, our study aims to develop a more comprehensive benchmark that rigorously evaluates spatial reasoning capabilities.
B.2 Spatial Reasoning in Psychology
In psychology, spatial reasoning refers to an individual’s ability to acquire, organize, utilize, and adapt spatial knowledge, recognized as one of the nine primary reasonings (Gardner, 2011). To systematically characterize this ability, Buckley et al. (2018)proposed a factor analysis-based framework, distinguishing between spatial visualization, spatial relations (such as mental rotation), and spatial orientation. Malanchini et al. (2020) further identified a strong correlation between spatial orientation and object manipulation skills. Additionally, Newcombe & Shipley (2014)introduced a classification system for spatial thinking, dividing spatial reasoning into two dimensions: intrinsic-extrinsic and static-dynamic. Furthermore, many studies also have contributed to the development of frameworks (Wai et al., 2009, Lee & Bednarz, 2012, Malanchini et al., 2020, Hegarty, 2010, McGarvey et al., 2018) and evaluation methods (Eliot & Smith, 1983, Uttal et al., 2024; 2013, Hegarty & Waller, 2005) for spatial reasoning. Their framework highlights the distinction between object-centric spatial properties and external reference frames, offering valuable insights for applications such as navigation and path planning. These psychological frameworks provide complementary perspectives for understanding and assessing spatial reasoning, offering theoretical foundations for embodied intelligent systems. Inspired by these classifications, our study proposes a novel taxonomy for visual-spatial reasoning, aiming to advance spatial reasoning research in vision-language models.
Table 7: Comparative analysis of various prompting and evaluation strategies.
Promopt Type
Eval Type Avg. Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
Manipulate
Motion
Analysis
Traffic
Analysis
Locate
Geospatial
Strategy
Pattern
Recognition
Geometric
Reasoning
Ego
Centric
Allo
Centric
Hypothetical
GPT-4.1-mini - - - - - - - - - - - -
None Direct 39.22 57.57 36.18 58.12 44.95 53.45 12.58 18.58 62.94 37.61 37.83
None RE 48.86 64.05 58.55 57.65 59.43 56.91 28.87 34.06 68.82 37.18 41.20
Zero-shot CoT RE 49.81 62.97 58.96 59.06 62.48 58.55 27.63 32.52 69.41 40.11 40.96
Manual CoT RE 49.76 65.68 58.90 58.59 64.38 56.91 28.45 32.13 69.61 39.31 41.20
None JSON 48.23 61.08 55.55 57.41 58.86 54.00 28.89 29.16 70.20 40.96 40.48
Zero-shot CoT JSON 48.70 58.67 57.17 57.88 57.71 55.09 28.89 29.29 69.02 40.16 42.89
Manual CoT JSON 48.87 64.32 56.53 59.06 60.19 56.36 29.28 30.19 72.55 39.57 39.28
None LLM 48.02 60.54 56.82 58.59 58.10 57.45 28.04 32.90 68.24 37.55 38.31
Zero-shot CoT LLM 48.36 64.05 56.71 58.59 58.86 57.27 27.84 33.29 67.06 38.30 38.80
Manual CoT LLM 49.85 62.97 59.48 58.12 61.33 58.36 28.04 31.61 69.02 41.12 39.28
Gemini-2.5-flash - - - - - - - - - - - -
None LLM 51.47 66.22 65.90 63.53 71.43 66.36 32.99 34.84 70.59 31.91 38.55
Zero-shot CoT LLM 51.53 63.51 61.27 58.82 67.62 65.45 42.27 34.84 79.41 35.90 32.53
Manual CoT LLM 52.12 67.57 62.72 68.24 73.33 60.91 38.14 34.19 75.49 35.90 33.73
Qwen-VL2.5-3B - - - - - - - - - - - -
None Direct 44.04 60.27 52.20 47.53 50.48 52.73 22.68 30.32 65.49 36.49 30.84
None RE 41.45 58.65 43.06 39.53 50.67 48.73 32.78 22.58 61.96 37.66 37.35
Zero-shot CoT RE 40.64 59.73 43.87 46.12 48.38 43.27 25.36 22.84 59.61 36.54 37.59
Manual CoT RE 40.07 55.68 46.65 47.29 40.57 46.00 28.04 24.39 60.39 33.35 31.57
None JSON 38.08 62.97 32.49 46.59 48.95 47.64 24.54 23.61 62.16 35.85 27.47
Zero-shot CoT JSON 39.20 61.35 40.40 48.94 45.52 47.27 22.27 24.00 65.29 33.51 27.71
Manual CoT JSON 38.37 58.11 34.28 42.35 46.29 48.55 27.63 24.65 62.75 36.54 26.75
None LLM 42.10 66.22 43.99 44.00 51.24 50.73 27.84 25.42 70.00 35.90 29.40
Zero-shot CoT LLM 35.89 54.59 36.24 49.41 40.19 40.18 24.95 20.13 61.37 29.20 33.98
Manual CoT LLM 36.26 53.78 34.10 50.35 42.10 44.00 23.51 26.06 60.39 31.38 23.86
Appendix C Additional Ablation Studies
C.1 Evaluation and Format Strategy
The choice of evaluation format and answer-extraction strategy can materially influence how fully a large language model’s capabilities are expressed. Because our benchmark introduces a novel visual–spatial reasoning task, Table˜7 reports the performance impact of different prompting and evaluation schemes. We consider three prompting paradigms—Direct QA, Zero-shot CoT, and Manual CoT—and three evaluation methods: (i) regex-based field extraction (RE), in which the VLM is instructed to place its answer in a fixed slot at the end of the response; (ii) JSON extraction, where the answer must appear in a dedicated answer field; and (iii) LLM-based free-form evaluation, for which we employ GPT-4.1-mini as an adjudicator.
We observe that reasoning-centric models that strongly obey formatting directives (e.g., Gemini-2.5-flash) cannot be reliably scored with regex extraction, because their compulsory chain-of-thought interferes with the required output template. For standard models, all three evaluation strategies yield comparable scores, and both CoT variants consistently outperform the Direct QA baseline by a small margin. Accordingly, in the main results table, we evaluate non-reasoning models with the Manual CoT + RE setting, whereas reasoning-oriented models are assessed with Manual CoT + LLM evaluation.
C.2 Evaluation on the full OmniSpatial benchmark
To complement the main experiments, we further conduct evaluation on the entire 8.4K OmniSpatial test set without applying the train/test split. This setting measures model performance on all available annotated questions, thereby reducing variance due to subset sampling and providing a more holistic view of spatial reasoning ability. Table 8 reports the results across all twelve fine-grained tracks.
Overall, the relative ranking of models remains stable compared to the smaller-scale splits discussed in the main text. Gemini-2.5-Pro achieves the highest average score of 55.05, securing Rank 1 among all compared systems. ChatGPT o3 follows closely with 54.52, while Gemini-2.5-Flash achieves 51.80. Other proprietary systems, such as GPT-4.1 and GPT-4o, also maintain consistent performance. On the open-source side, Qwen-VL-2.5-32B achieves the strongest result with 43.97, outperforming other Qwen variants and aligning with the trends observed in the evaluations.
A closer per-track analysis highlights complementary strengths across models. ChatGPT o3 excels at Manipulate, Traffic Analysis, and also leads in Pattern Recognition, Ego-Centric, and Hypothetical reasoning. In contrast, Gemini-2.5-Pro dominates Locate, Geospatial Strategy, Geometric Reasoning, and Allo-Centric perspective taking. Gemini-2.5-Flash yields the highest accuracy on Motion Analysis. These complementary strengths suggest that different architectures capture distinct aspects of spatial reasoning, and point towards potential benefits of ensembling or multi-expert distillation.
Table 8: Evaluation on the full OmniSpatial benchmark. Dark green marks the best Avg. accuracy and light green marks the second-best Avg. accuracy in all the models.
Method
Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
Manipulate
Motion
Analysis
Traffic
Analysis
Locate
Geospatial
Strategy
Pattern
Recognition
Geometric
Reasoning
Ego
Centric
Allo
Centric
Hypothetical
Proprietary Models
GPT-4o 44.47 5 60.00 57.97 62.38 55.24 59.09 29.90 31.01 42.06 24.34 28.05
GPT-4.1 48.75 4 61.18 66.67 66.60 69.52 60.91 30.93 35.58 47.06 27.26 33.06
o3-04-16 54.52 2 71.05 63.64 70.61 69.05 61.70 42.55 47.15 53.35 30.98 46.28
Gemini-2.0-Flash 42.41 7 56.47 56.76 55.95 63.81 57.27 15.46 30.04 43.20 24.56 28.05
Gemini-2.5-Flash 51.80 3 54.84 70.43 65.33 66.00 68.75 39.53 39.46 52.70 33.49 36.84
Gemini-2.5-Pro 55.05 1 66.67 68.34 69.10 78.85 71.15 36.17 50.17 50.65 39.57 37.89
Open-source Models
Qwen-VL-2.5-3B 39.76 8 57.65 47.34 49.07 49.52 58.18 29.90 26.97 40.62 30.47 29.88
Qwen-VL-2.5-32B 43.97 6 58.82 52.42 52.97 68.57 51.82 27.84 29.51 51.70 26.82 35.98
C.3 Details of Human Baseline & Inter-Annotator Agreement
To establish an upper bound of performance, we conducted a human evaluation on the OmniSpatial benchmark. Six human annotators (graduate students with backgrounds in vision and robotics) were recruited. Each participant was presented with a randomized subset of questions covering all four tracks and a balanced selection of fine-grained tasks. The interface was blinded: multiple-choice options were randomized and no additional context was provided. Annotators were instructed to answer independently without external resources.
Each item was labeled by three different annotators, and the final answer was obtained via majority vote. To quantify annotation reliability, we report inter-annotator agreement (IAA) using Krippendorff’s $\alpha$ for multi-annotator agreement. Table 9 summarizes the results across tracks. We also report 95% confidence intervals (bootstrap over questions) for human accuracy.
Table 9: Human baseline accuracy and inter-annotator agreement (IAA) across tracks.
Track Accuracy (%) Krippendorff’s $\alpha$
Dynamic Reasoning 95.2 $\pm$ 1.3 0.92
Spatial Interaction 93.5 $\pm$ 1.6 0.85
Complex Logic 87.9 $\pm$ 5.5 0.76
Perspective Taking 94.4 $\pm$ 2.2 0.80
Overall 92.6 $\pm$ 1.5 0.84
These results indicate that human annotators achieve $\sim$89% average accuracy, with substantial agreement across annotators ($\kappa$$\alpha\approx 0.84$). This provides a reliable estimate of the human upper bound and confirms the internal consistency of the benchmark.
Appendix D System Prompts
We present all the system prompts used in our experiments in Figs.˜7 and 8 to facilitate reproducibility. We observe that some models are sensitive to the choice of system prompt, which may stem from distributional biases in their training data. In contrast, inference-oriented models generally exhibit stronger generalization capabilities and are less reliant on carefully crafted prompts.
Appendix E Additional Visualization
We present more examples of question-answer pairs from the OmniSpatial, with perspective taking, spatial interaction, dynamic reasoning, and complex logic, shown in Figs.˜9, 10, 11 and 12, respectively. Our benchmark comprises a rich collection of data samples spanning diverse scenarios, resolutions, lighting conditions, and geographical regions. It includes both absolute numerical analyses and relative spatial relationships, aiming to comprehensively evaluate the spatial reasoning capabilities of vision-language models.
Appendix F Limitation & Future Works
Although OmniSpatial includes some image clips with dynamic information from HOI4D (Liu et al., 2022), the complexity of the operational tasks still lags behind that of long videos (Yang et al., 2024, Chandrasegaran et al., 2024). Moreover, while PointGraph & Spatial CoT enhances VLM’s spatial understanding through point cues, the improvement is not fundamental in nature. Spatial reasoning tasks are more like mathematics and coding tasks, require longer and more complex reasoning (DeepSeek-AI et al., 2025, Jaech et al., 2024).
3D information is crucial for spatial reasoning, and future work involves introducing 3D representation (Dong et al., 2023, Qi et al., 2023a; 2024; b) and perception (Qi et al., 2017a; b, Zhang et al., 2024a, Fan et al., 2024), as well as 3D VLMs (Peng et al., 2024, Dong et al., 2024, Guo et al., 2023, Qi et al., 2024; 2025), reasoning model (Zhang et al., 2025a, DeepSeek-AI et al., 2025, OpenAI, 2025b) and knowledge distillation (Zhang et al., 2023, Hinton et al., 2015, Dong et al., 2023, Qi et al., 2023a). The ultimate goal of spatial reasoning is to empower robots, and future work also involves robot execution tasks (Huang et al., 2024a, Fang et al., 2024a, Qi et al., 2025, Huang et al., 2024b, He et al., 2025b).
Appendix G Broader Impacts
OmniSpatial promises several positive societal benefits. By pushing models to reason about motion, collision risk and traffic scenes, it can hasten the arrival of safer autonomous vehicles and service robots that foresee hazards and navigate crowded spaces responsibly. Its geometric-reasoning tasks—from polyhedron unfolding to assembly—offer data that can streamline product design, packaging and manufacturing, lowering material use and energy waste. The benchmark’s localization, UI-interaction and perspective-taking challenges cultivate spatially aware assistants that improve AR/VR experiences, access tools for visually impaired users and more natural human-computer interfaces.
Figure 7: System prompts used in OmniSpatial evaluation.
Figure 8: System prompts used in OmniSpatial evaluation.
Figure 9: Visualization example of OmniSpatial data samples.
Figure 10: Visualization example of OmniSpatial data samples.
Figure 11: Visualization example of OmniSpatial data samples.
Figure 12: Visualization example of OmniSpatial data samples.